Identify the correct graph of the following piecewise-defined function. -3x - 6 ifx-2 if x > -2 8(x) = Answer 2 Points
Answer #1

w -32-6 24-2 Arya foot u 23-2 al 452 for - 3x-6 at x = -2 foo2, 6-2= y (-2,42 6-2012 ad X:-2 at xa-2 FO()= 4 4 - 2 Any